art A A 72-gram sample of natural carbon contains the usual fraction of 14C which is 1.3×10-10%. Estimate roughly how long it will take before there is only one C nucleus left. Express your answer using two significant figures. t= 2.4x 105 yr


given this info, find

given this info, find
Answer #1

PARTA

one 14/ 6C nucleus = 6 * 1.6726 * 10-27+ 8 * 1.6749 * 10-27 =    23.4348 * 10-24gm

half life of   14/ 6C =   5730 years

decay constant = ln2/ 573 =   1.2096 * 10-4   yr-1

23.4348 * 10-24 = 9.36 * 10-11 * e(-1.2096 * 10-4 * t)

t = 239879.542   years

PART B

if the sample is 340 grams

23.4348 * 10-24 = 4.42 * 10-10 * e(-1.2096 * 10-4 * t)

t = 252712.54 years